A stock just paid a dividend of $1.05. The dividend is expected to grow at 23.56% for three years and then grow at 3.43% thereafter. The required return on the stock is 10.68%. What is the value of the stock? Answer Format: Currency: Round to: 2 decimal places.
Answer #1

D1=(1.05*1.2356)=1.29738

D2=(1.29738*1.2356)=1.603042728

D3=(1.603042728*1.2356)=1.980719595

Value after year 3=(D3*Growth Rate)/(Required return-Growth Rate)

=(1.980719595*1.0343)/(0.1068-0.0343)

=$28.25735555

Hence current value=Future dividends*Present value of discounting factor(rate%,time period)

=1.29738/1.1068+1.603042728/1.1068^2+1.980719595/1.1068^3+$28.25735555/1.1068^3

=$24.78(Approx).